What is the weight of a steel plate 15 ft. long, 4 ft. wide, and 5/8 inch thick, if the steel weights 475 lb./ft.^3?


1
Expert's answer
2021-06-28T03:40:32-0400

the formula of to find weight is "m = \rho V = \rho abc"

"a = 15ft", "b = 4ft", 1 inch is equal to  "\frac{1}{12} ft:" "c = \frac{1}{12}*\frac{5}{8}=\frac{5}{96}ft"

"\rho = 475 \frac{lb}{ft^3}"

"m = 15ft*4ft*\frac{5}{96}ft*475\frac{lb}{ft^3}=1,484.375lb"